1. Why was a transverse sensitivity correction necessary for the transverse strain gage and not the longitudinal?

2. How could Poisson's Ratio be measured on a cylindrical concrete specimen subjected to axial compression?
Answer #1

1) Transverse sensitivity correction is necessary for the transverse strain gage because

    In plane wire strain gages , transmission of strain in to the wire from a direction perpendicular to the wire axis is nearly negligible .As a result , the transverse sensitivity of these gages is due almost entirely to the fact that a portion of the wire in the end loop lies in the transverse direction.

2) For a cylindrical concrete specimen subjected to axial compression ,

     the length of the cylindrical specimen will be reduced and the diameter will be increased.

      Let us assume the initial length be L

                               initial diameter d

      After the application of compressive force which causes a compressive stress

                  the length is reduced to l

                and the diameter is increased to D

         change in length = L -l = \Delta L

       change in diameter = D - d = \Delta D

\therefore linear strain = change in length / original length = \Delta L / L

     lateral strain = change in diameter / original diameter = \Delta D / d

Poissons Ratio = lateral strain / Linear strain = (\DeltaL / L) / ( \Delta D/d )